OpenStudy (anonymous):

The number of members in a club dropped from 80 to 52. What is the percent of decrease in the number of members?

OpenStudy (anonymous):

I NEED HELP PLEASE

OpenStudy (anonymous):

ok step by step or just the answer?

OpenStudy (anonymous):

the answer

OpenStudy (anonymous):

So first step we want to know how much decrease there already is IN NUMBERS so to find that out we use the equation x = amount of decrease 80 - 52 = x 80 - 52 = 28 x = 28 So now we know that 28 people left the club Next step, we want to know the PERCENT of decrease So we divide the AMOUNT of decrease by THE ORIGINAL NUMBER OF MEMBERS 28/80 = 0.35 0.35 is a decimal, but it can be made into a percent by multiplying it by 100 0.35 * 100 = 35 So there was a 35% percent decrease YOUR ANSWER: 35%

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

do you have any more questions? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok post them I'll try to answer

OpenStudy (anonymous):

The price of a bracelet increased from $170 to $200 on Mother's Day What is the percent of increase in the price? Round to the nearest percent.

OpenStudy (anonymous):

Answer

OpenStudy (anonymous):

200 - 170 = 30 30/200 = 0.15 0.15 * 100 = 15 The percent of increase in the price is 15%

OpenStudy (anonymous):

Thanks that is all

OpenStudy (anonymous):

yw :)